OpenStudy (anonymous):

HELP PLEASE what are the zeros for the function f(x) = 2x2 + 8x + 10 and how many times does the graph cross the x-axis?

OpenStudy (anonymous):

@Rushwr

OpenStudy (johnweldon1993):

Do you know how to factor quadratics? or use the quadratic formula?

OpenStudy (anonymous):

i got x=-2+-i

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

yes that is correct. however that also means that the quadratic never crosses the x axis as it only has complex roots

OpenStudy (anonymous):

Thank you so so much :))) can you help with a few more? if not its okay!

OpenStudy (anonymous):

in this case I should say.

OpenStudy (johnweldon1993):

Im sure someone can help @Morgan27 Go ahead and post it :)

OpenStudy (anonymous):

If a parabola's focus is at (−2, 4) and the directrix is at y = −2, what is the vertex form of the equation representing this parabola?

OpenStudy (johnweldon1993):

Oh I'm glad I checked back in oops...still need help with this one?

OpenStudy (anonymous):

yesss i do lol thank you

OpenStudy (johnweldon1993):

Well just in case... The formula we will be using is \[\large (x - h)^2 = 4p(y - k)\] first thing we need to find is the vertex (h,k) The vertex is halfway between the focus and the directrix....so halfway between y = 4 and y = -2 would be at y = 1 So we know the vertex is located at (-2,1) Now we just need the distance from the vertex to the focus...which is 3 (p) Now we have everything we need \[\large (x - (-2))^2 = 4(3)(y - 1)\] \[\large (x + 2)^2 = 12(y - 1)\] Now we just solve for 'y' \[\large y = \frac{1}{12}(x + 2)^2 + 1\]

OpenStudy (anonymous):

y=4/3?

OpenStudy (johnweldon1993):

Huh? We just needed the vertex form of the equation :)

OpenStudy (anonymous):

ohhhh lol im sorry

OpenStudy (johnweldon1993):

No problem at all :)

OpenStudy (anonymous):

\[y=(x+2)^2+\frac{ 1 }{ 12 }+1\]

OpenStudy (johnweldon1993):

Nope...the 1/12 is multiplied to the (x+2)^2

OpenStudy (anonymous):

i got it now. Thank you for all of your help! :)

OpenStudy (johnweldon1993):

Sorry for the REALLY late reply but you're welcome :)